Solve the equations by finding the exact solution.

4 = x − 4

Answer by jim_thompson5910(35256)

Add 4 to both sides to get

4 = x - 4

4 + 4 = x - 4 + 4

8 = x + 0 ... Notice how the -4 and +4 add to 0

8 = x

x = 8

-------------------------------------------------------

Final Answer: x = 8